OpenStudy (anonymous):

Which literal equation is not equivalent to x equals two y plus z ? x minus two times y equals z x minus z equals two times y y equals x minus z divided by two x plus two y equals z

OpenStudy (anonymous):

x = 2y + z x - 2y = z (correct) x - z = 2y (correct) y = (x - z)/2 (correct) x + 2y = z (incorrect) <-- your answer